OpenStudy (anonymous):

PLLLEEZZZ! solve: 7^(x-5)=7^(4x+7)

OpenStudy (australopithecus):

take log of base 7 on both sides to get x-5 = 4x + 7

OpenStudy (anonymous):

\[\large x^a = x^b \implies a = b\]

OpenStudy (anonymous):

No need to take log here..

OpenStudy (anonymous):

Just equate the exponents...

OpenStudy (anonymous):

i dont understand...

OpenStudy (anonymous):

since bases are same so just equate both sides \[\Large (x-5)=4x+7\] take all the x terms to one side \[\Large 4x-x=-12\] \[\Large 3x=-12\] i hope you got it.

OpenStudy (anonymous):

Tell me what are the exponents both the sides ??

OpenStudy (anonymous):

@cassie4math did you get it?

OpenStudy (anonymous):

I give you the example: \[\huge 5^{x-6} = 5^{-2x}\] This implies : \[\large x - 6 = -2x\]

OpenStudy (anonymous):

so x=-4 riught?

OpenStudy (anonymous):

Well Done..

OpenStudy (anonymous):

thanks!

OpenStudy (anonymous):

Welcome dear..